OpenStudy (perl):

A box contains 8 dark chocolates, 8 white chocolates, and 8 milk chocolates. I choose chocolates at random (yes, without replacement; I’m eating them). What is the chance that I have chosen 20 chocolates and still haven’t got all the dark ones? what is your answer to that

OpenStudy (anonymous):

dont really get it

OpenStudy (usukidoll):

choose milk chocolates choose white chocolates and just 4 dark chocolates XD

OpenStudy (anonymous):

the question says that you still haven't got ALL the dark ones, so you COULD get 4 dark chocolates to 7 dark chocolates, but not the full 8

OpenStudy (perl):

oh woops

OpenStudy (perl):

i thought it said 'any'

OpenStudy (anonymous):

do you get the question now?

OpenStudy (perl):

yes

OpenStudy (usukidoll):

I'm thinking milk white milk white milk white milk white milk white

OpenStudy (perl):

you wantthe probability of not getting all 8 dark chocolates. you are guaranteed at least 4 dark chocolates so the possibilities are 4 dark, 5 dark, 6 dark, 7 dark

OpenStudy (anonymous):

yep, so if you find the probability of getting 8 dark chocolates, you can minus it from 1 and you'll get the probability of getting 4 dark, ... 7 dark

OpenStudy (perl):

there are 4 favorable cases, out of 5 total possible cases favorable cases : : exactly 4 dark , 16 non-dark : exactly 5 dark 15 non-dark exactly 6 dark 14 non-dark exactly 7 dark 13 non-dark total cases : exactly 4 dark , 16 non-dark : exactly 5 dark 15 non-dark exactly 6 dark 14 non-dark exactly 7 dark 13 non-dark exactly 8 dark , 12 non-dark I see no reason not to treat these choices as equally likely. therefore the probability of not picking all dark ones is 4/5

OpenStudy (perl):

since they must add up to 20 , these are the only possibilites

OpenStudy (perl):

notice 3 dark 17 non-dark is impossible

OpenStudy (perl):

the situation is constrained by two facts there must be no more than 16 non-dark chocolates there must be at least 4 dark chocalates

OpenStudy (perl):

that produces (16,4) (15, 5 ) (14,6) (13, 7) (12,8)

OpenStudy (perl):

the situation is constrained by two facts you cannot have more than 16 non-dark chocolates you cannot have more than 8 dark chocolates there must be at least 4 dark chocalates so the possibilities are (# non-dark, # dark) (16,4) (15, 5 ) (14,6) (13, 7) (12,8) there is 4/5 favorable

OpenStudy (perl):

can i get a yes?

OpenStudy (anonymous):

sure, i guess

OpenStudy (perl):

lol

OpenStudy (anonymous):

the explanation seems reasonable enough

OpenStudy (perl):

i think its more complicated, let X = number of dark ones , Y = number of non-dark chocolate how many ways can you choose 16 non dark and 4 dark? how many ways can you choose 15 non-dark and 5 dark?

OpenStudy (perl):

P( chance not getting all dark in 20 draws) = P( X = 4 & Y= 16 or X= 5 & Y=15 or X=6 & Y=14 or X=7 or Y=13 )

OpenStudy (anonymous):

that's what i thought too, but if you add up all the ways to not get all 8 dark chocolate, i'm like 90% sure you'll get 4/5...cause there are like 60 ways if you account for each individual selection

OpenStudy (perl):

P( X = 4 & Y= 16 or X= 5 & Y=15 or X=6 & Y=14 or X=7 or Y=13 ) = P ( X=4 & Y=16) + P(X=5 & Y=15) + ...

OpenStudy (anonymous):

btw, what level of maths is this?

OpenStudy (perl):

not sure

OpenStudy (anonymous):

i meant like are you still in school or university?

OpenStudy (perl):

P(Y=16 & X=4) = P(Y=16) * P (X=4 | Y = 16) =

OpenStudy (perl):

im in university

OpenStudy (anonymous):

oh, then you might not want to listen to me, i'm in yr 11 doing yr 12 maths so yea...

OpenStudy (perl):

i find this problem confusing

OpenStudy (anonymous):

whoops, should have told you that at the start!

OpenStudy (perl):

im trying to force these into mutually exclusive possibilities

OpenStudy (perl):

you cannot have more than 16 non-dark chocolates you cannot have more than 8 dark chocolates there must be at least 4 dark chocalates so the possibilities are (# non-dark, # dark) (16,4) (15, 5 ) (14,6) (13, 7) (12,8)

OpenStudy (perl):

but are these equally likely?

OpenStudy (perl):

if these possibilities are equally likely to occur (and they are mutually exclusive), then the probability is 4/5

OpenStudy (anonymous):

i'd assume so, unless some were heavier or bigger than the others (which isn't so)

OpenStudy (perl):

im use to working on problems like 5/12*4/11 , those sorts of problems

OpenStudy (anonymous):

me too, I guess you could do that, but you'd be there forever!

OpenStudy (perl):

what was your approach?

OpenStudy (anonymous):

I found the probability of getting all 8 chocolates, kinda took a while and subtracted it from 1

OpenStudy (perl):

can you show me your work

OpenStudy (perl):

i need your help :)

OpenStudy (perl):